UT Dec 09 - Split Night LBTB/LN-COM and Eng (MODS Restart)



Night Info

  • Sunset 17:14
  • Nautical Twilight Ends/Moonrise 17:41
  • Astro Twilight Ends 18:11
  • Night Center 00:12
  • Handover from LBTB/LN to LBTO/MODS Eng 00:07 (split the reconfig from LN to MODS)
  • Astro Twilight Starts 05:43
  • Nautical Twilight Starts 06:12
  • Sunrise 07:10
  • Moon illumination ~99%

Staff

  • OSA: S.Allanson
  • ISp: E.Solheid
  • Mtn Manager: B.Bingham
  • ISA/SO: O.Kuhn

Night Plan

  • LBTB (either LN-Com or Science) first half

MODS Restart On-sky Checkout

  • will check times - but approximately 2.5-3 hours. 00:18 - 03:15 MST
  • AGw checkout (~1 hr)
    • Check WFS hotspot
    • Check alignment of rotator center on guider to WFS hotspot
  • Rotator Center on MODS (15-20 min): M52, Hyades, Coma Star Cluster (use pointing mask)
  • Rotator Zero Point (try for 2 masks, 15-min per mask): masks for 00h25d, 04h25d,06h80d and 08h25d to cover night
  • From 20220925 wiki: Rotational offsets needed to align MODS MOS masks (2021B to 2022A):
instr mean median sig_n-1 n
MODS1 -0.03 -0.03 0.05 33
MODS2 -0.05 -0.04 0.05 55
  • Spectrophotometric Standard (roughly 30-min): 1spec/mode mode if non-phot, 3 spec/mode if phot
    • All stds are accessible but G71, G191-B2B are available most of the night. G191-B2B will take about 45-min (Nimgs=3) or 15-min (Nimgs=1) not counting acquisition.
  • long-slit alignment and blind offset check on mask field with 2.4" slit (20-min)
  • Dithering Imaging Scripts (10-min): on mask field with radec and detxy offsets (this is to see if we run into any offset errors)

03:15-03:31 MST (16-min reconfig time)

Rollover of TMS/LBC Work

03:31-06:12 MST
  • Fields

12-deg twilight at 06:12 MST

Night Log

Summary:

We could not open all night. The LN team worked during the first half and we remained configured for about an hour after handover to allow them to continue, since the humidity was pegged at 99.9%. At 2:30, Heejoo and Yang made some tests of the TMS channel-dropping algorithm, and these seemed successful although Heejoo will make a more thorough analysis. The humidity went down, but clouds moved in, so the TMS testing continued. With MODS, I obtained sets of biases and a good set of calibrations (blue-only pix flats need to be finished) and a dual-grating focus sequence. IMCS TTF zeropoints were measured and implemented.

Log:

22:30-22:40 Around 15:30 or 15:40 MST MODS were filled with LN2. The coldStart scripts were run, after which Chris & Elliott loaded the masks. I updated the mask tables but did not run the maskSnap yet.

I could not power up the MODS1 Blue and MODS2 Red HEBs initially. Tried a few times and Chris & Elliott reset the breakers. Eventually I could remotely power up the MODS1 Blue HEB, but the MODS2 Red HEB had to be set to bypass the remote toggle. I filed IT 8782 (High Priority).

00:17 Staying closed at sunset due to clouds/fog/humidity.

00:46 Ran the sieveSnap

00:50 Ran the IMCSTTF script at 60-deg elevation & LDG/RDG 207/195 deg.

01:11 The CCD temperatures are just now within nominal range - only MODS2R is in the 'red' range. 01:14 it went into the nominal range.

01:15 Elliott is topping off the dewars with LN2

02:40 I updated the [b|r]config.tab files on mods1 and mods2 and closed/reopened the GUIs. The IMCS TTF zeropoints changed very little, in some cases < 5 mic, for MODS1, but the changes were larger, ~100-200 mic, for MODS2. I think that this is what I noted on restart also so will look at this more (the changes now are greater than they were at restart). IMCS locked on first try during the running of the script.

02:47 I started the MODS alarm service following Yang's instructions (this is a new script, replacing the old emailMODS[2]errors.sh scripts).

04:50 Taking 50 LBC biases on each side. Dome is dark.

06:04 MODS biases: 15x 8Kx3K, 3Kx3K and 1Kx1K biases. Notes:
  • Noticed m2b image counter is not advancing, although series of images are being taken (advanced from 1 to 2 at 06:15 UT but after a few biases were taken).
  • The M1b nextfile name was not fully filled in (update fixed this) at 06:56.
  • MODS2B 100-sec delays are pretty frequent, but started only during the 3K & 1K series. Between 1/2 and 1/3 of these images have a delay. MODS1R 100-sec delays are happening also but at a lower frequency and for all ROIs.
  • blue and red fitsflush on MODS2 at 00:13 to get the last few 1K biases.
ROI m1b m1r m2b m2r
8Kx3K 13-27 12-26 13-28 13-27
3Kx3K 28-42 27-41 29-43 28-42
1Kx1K 43-57 42-55 44-58 43-57
07:10 Handover was 00:07, although RH is still 99.9% and now we have clouds. We'll remain at zenith, with LN configured to allow them to continue while I continue taking MODS calibrations.

07:14 more MODS calibrations
Calib m1b m1r m2b m2r
dual grating lamps 58-60 57-59 59-61 58-60
dual grating pix flats 61-70 60-64 62-71 61-65
dual imaging flats ugriz 71-80 65-79 72-81 66-80
red imaging flats riz   80-94   81-95
blue imaging flats ug 81-90   82-91  
red grating lamps   95-97   96-98
red grating pix flats   98-   99-
blue grating lamps 91-93   92-94  
blue grating pix flats 94-98 (unfinished)   95-98 (unfinished)  
08:11 Steve is reconfiguring for MODS. (If we still cannot open, we'll reconfigure to the LBCs to run a few TMS closed-dome tests) and then go back to MODS.

09:31 Steve reconfigured to the LBCs for closed-dome tests. Heejoo and Yang are on the polycom for closed-dome TMS testing.
  1. Channel signal strength on the Sx side (since VxWorks issue last night)
    1. checked - on SX it is ok, but on DX gains needed to be increased. We took 2 references, both had the same channels as a few nights ago (20221202)
  2. python code function test for wiki update
    1. ./tms_loop.py -p does not run in passive mode; corrections are being made. wiki page needs to be updated & code needs to be worked on.
  3. Dropping channels: ./tms_drop.py
    1. Zenith:
      • 10:55 (just before): Dropped channels 15 and 22: -15, -22
      • 11:03: Dropped channel 10 in addition to 15 & 22. SX corrections are varying a bit, more than DX
      • 11:10: Dropped channel 4 in addition to 15, 22 and 10. Running TMS with minimum required channels.
      • 11:18: Enabled all channels (in prep for moving to another elevation) 2 60-deg:
      • 11:26 At 60-deg, all channels enabled.
      • 11:33 Dropped 15 and 22
      • 11:39 Dropped 10 in addition to 15, 22.
      • 11:41 Dropped 4 in addition to 10,15 and 22.
      • 11:46 Enabled all channels. Then we'll try the 1, 1, 4 combination (4 on one retro, 1 on each of the other 2).
      • 11:49 Dropped channels 5, 6 and 11.
      • 11:53 Moved SX mirror in Y by 0.5 mm with channels 5, 6 & 11 dropped...(wrong side...channels 5,6,11 are on DX)
      • 11:55 Moved DX mirror in Y by 0.5 mm with channels 5, 6 and 11 dropped. TMS did not move the mirror back, though. Not enough valid channels.
      • 12:02 after enabling all, finally TMS moved the mirror back. 2 30-deg:
      • 12:09 Stabilized at elev 30-deg, though there are some channels being filtered out by stddev filtering 2 40-deg:
      • 12:16 Stabilized. -15, -22
      • 12:21 Dropped channel 10 in addition to 15, 22. Now both sides have the minimum number of channels (6). 17,20,21,23,24,26 on sx and 4,5,6,11,12,13 on dx.
      • 12:26 enall and allow it to stabilize
      • 12:28 Dropped channels 5 and 10 (2,2,2 combination on DX).
      • 12:41 enall -- all channels back except 2 which is missing at this elevation.
      • 12:43 sending an OB which changes the filters to U-Bessel and r-SLOAN. There were no filter focus offsets in our earlier tests, since the LBCs were not being used.
      • 12:48 Dropping channels 15 and 22 with U-Bessel & r-SLOAN.
      • 12:52 Changing filters to V-Bessel & z-SLOAN
      • 12:59 Stop ./tms_loop.py --mcp both and add the -o option: ./tms_loop.py --mcp both -o 2 13:04 Slew to track at RA/DEC = 10:00:00/20:00:00 70 deg
      • Channel 2 came back
      • 13:09 Dropping channels 15, 22
      • 13:15 Dropping channel 10 in addition to 15,22.
      • 13:22 Dropping channel 4: Now 4,10,15,22 all dropped.
  4. measurable off-pointing distance of the collimator: translate mirror and see std filtered out channels.
    • zenith with TMS corrections applied there.
    • Run tms_loop_passive.pl as we adjust DX M1 Y position.
      • Total collimation was -1.120, added +0.5mm at 13:38 --> total collimation -0.620, no filtered.
      • Another +0.5mm step at 13:40 --> total Y collimation -0.120
      • Another +0.5mm step at 13:43 --> total Y coll 0.380 mm. All channels came back.
      • Another +0.5mm step at 13:45 --> total Y coll 0.880. All channels came back.
      • Another +0.5mm step at 13:46 --> total Y coll 1.381. Lost only channel 2.
      • Another +0.5mm step at 13:49 --> total Y coll 1.88. Now 2 and 6 are missing.
      • Another +0.5mm step at 13:51 --> total Y coll 2.38. Still just 2 and 6 missing.
      • Another +0.2mm step at 13:54 --> total Y coll 2.58. Still just 2 and 6 missing.
      • Now the other way, a -5mm step at 14:00 --> total Y coll -2.419mm. Now channels 2 and 5 are missing.

11:36 MODS dual grating focus run launched. During TMS testing, while at 60 deg elevation.

script m1b m1r m2b m2r Comments
dual grating focus -108 -112 -109 -114 M2 finished, M1 did not before slew to 30-deg
14:04 TMS test ended. 14:09 LBCs are off. 14:14 Putting MODS to sleep (need to try to get blue-only pix flats at some point).

Over to the day crew for instrument fills. Elliott will toggle the MODS2R back to remote and later today we'll see if toggling the switch back and forth will help.

-- BarryRothberg - 06 Dec 2022
Topic revision: r11 - 09 Dec 2022, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback